OpenStudy (anonymous):

Can anyone help me with this: Evaluate the formula A = TT r 2 for TT = 22 over 7 and r = 7 over 4

OpenStudy (wolf1728):

Area = PI • r² where PI = 22/7 r = 7/4 Area = (22/7) • (7/4) • (7/4) Area = (22) • (1/4) • (7/4) Area = 9.625

OpenStudy (anonymous):

Thanks Wolf..you rock.

OpenStudy (wolf1728):

Thank you :-)